# Evaluating Foundation Models’ 3D Understanding Through Multi-View Correspondence Analysis

### v1

@@ 4.1.2 Discussion
Although none of the models were explicitly trained for viewpoint understanding, the results show clear differences in their ability to generalize across views. As seen in the previous section, DINOv3 performs best, followed by DINOv2 and DINO. This indicates that the features of the DINO models, compared to other models, capture object shape more reliably and remain consistent under viewpoint changes. This consistency may arise from their self-supervised training objective, which aligns features across different augmentations of the same image and promotes stable relations between local and global cues, leading to structured representations where similar object parts share similar embeddings. DINOv3 further benefits from large-scale pretraining and the Gram Anchoring mechanism [DINOv3], leading to more stable local features and stronger geometric consistency under viewpoint changes.
While DINO performed better under the easier difficulties, where multiple views of the object were available, DINOv2 obtained a higher mIoU under the Extreme difficulty with only one reference bin. This result suggests that DINO’s simpler self-distillation loss favors interpolation across views, whereas DINOv2’s larger-scale self-supervised pretraining supports stronger extrapolation to unseen angles. CLIP generally ranked third, reflecting some ability to encode visual-semantic regularities despite being trained for image-text alignment rather than geometric consistency. By comparison, SigLIP2, TIPS, and C-RADIOv2 performed substantially worse, indicating that objectives centered on semantic matching or mixed supervision may weaken part-level consistency needed for viewpoint-robust segmentation.
We theorize that the consistent underperformance of VGGT stems from a mismatch between its architecture and the Hummingbird retrieval setting. VGGT is designed as a geometry-grounded, multi-view transformer whose aggregator learns to combine information across several aligned frames of the same object or scene [VGGT]. The original VGGT paper does not include dense segmentation or retrieval-style evaluation; instead, it focuses on 3D geometry tasks such as camera pose, depth, and surface normal estimation. In our evaluation, the model receives only a single view ($S=1$), meaning the aggregator cannot make use of its multi-view fusion objective. Moreover, VGGT does not provide ViT-style patch embeddings, so we adapt its aggregator outputs to the patch-token interface required by Hummingbird. Together, these factors lead to the weak dense semantic segmentation performance observed.
Figure 2: Segmentation performance across viewpoint bins. Each subplot (a)–(d) corresponds to a different difficulty level. Bars show mIoU scores on unseen validation bins, while dots show performance on the reference bins stored in the memory bank. The figure illustrates how segmentation accuracy declines as the validation angles move farther from the reference viewpoints. DINOv3, DINOv2, DINO, and CLIP maintain higher, more stable performance as angular distance increases, whereas other models degrade more quickly. VGGT consistently underperforms all other models, yet still follows the expected degradation pattern observed across difficulty levels. Overall, the trend highlights how self-supervised pretraining leads to smoother cross-view generalization.
@@ 4.2 Experiment B: breaking point analysis
In this experiment, we analyze whether the models experience a sudden failure in viewpoint generalization, and if so, at what angle. The goal is to identify the angular range within which performance remains stable under the Extreme difficulty. In particular, we perform a breaking point analysis under the Extreme setting from experiment A, where only a single reference bin (0) is available in the memory and models are evaluated across the remaining bins. We define the breaking point as the earliest validation bin where a model’s normalized performance drops significantly compared to the previous bin. Specifically, we compute the performance drop $\Delta_{i}$ per bin $i$ as
$\Delta_{i}=$ $\text{ norm~mIoU}_{i}-\text{norm~mIoU}_{i-1},$ (2)
$\text{norm~mIoU}_{i}=\frac{\text{mIoU}_{i}}{\text{mIoU}_{0^{\circ}}},$ (3)
where the normalization by the 0 bin ensures that performance drops reflect viewpoint sensitivity rather than overall model scale.
@@ 4.2.1 Results
As illustrated in Figure 3, most models show a gradual decline in normalized mIoU as the validation bin angle increases. TIPS and VGGT are the models that reach a breaking point, both at the 30 bin (see Table 3). TIPS has a sharper decrease than C-RADIOv2, with a decrease of 0.115 between 15 and 30. VGGT’s performance drop is significantly larger: it drops by more than 0.2 between 15 and 45. In contrast, DINOv2, DINOv3, CLIP, DINO, and SigLIP2 degrade smoothly without reaching a breaking point.
Figure 3: Normalized mIoU under viewpoint shifts. Each curve shows the normalized mIoU of a model, measured relative to its 0° reference-bin performance. The horizontal axis represents increasing viewpoint angles (validation bins). A steep drop between consecutive bins corresponds to a potential breaking point in generalization. We can see DINOv2 and DINO degrade more smoothly than TIPS and C-RADIOv2, especially between 15 and 30. This may indicate that the latter models have weaker viewpoint robustness. DINOv3 peaks at the 60°validation bin, and then drops below the performance of DINOv2 and CLIP, resulting in an inconsistent degradation curve.
Table 3: Breaking points. For each model in experiment B, we report the breaking point validation bin where the normalized mIoU drop exceeds 10% ($\Delta_{i}\leq-0.1$) relative to the previous bin. “None” indicates gradual degradation without a sudden breakdown. Only TIPS and C-RADIOv2 reach a breaking point, both at 30°, whereas DINO, DINOv2, CLIP, and SigLIP2 maintain smoother decline curves.
Model Breaking point bin norm. mIoU drop
CLIP ViT-B/16 None
DINO ViT-B/16 None
DINOv2 ViT-B/14 None
DINOv3 ViT-B/16 None
C-RADIOv2 ViT-B/16-CPE None
SigLIP2 B/16-512 None
@@ 4.2.2 Discussion
The breaking point analysis highlights differences in robustness across models and shows how their pretraining objectives affect stability under viewpoint shifts. DINOv2 is the most stable, maintaining gradual degradation without sudden drops. CLIP shows a similar trend, with slightly weaker overall performance. DINOv3 follows DINOv2 and CLIP closely and is the only model that shows a slightly fluctuating degradation pattern. Potentially, this is due to architectural factors such as the influence of register tokens and its more complex pretraining objective. Alternatively, it may be affected by the reduced per-bin sample counts noted earlier. More research is needed to clarify the contribution of these factors. The normalized performances of DINO and SigLIP2 also decrease steadily, with a slight increase at the 90 bin. Interestingly, CLIP and SigLIP2 preserve some resilience despite being trained for image-text alignment rather than spatial structure. Although C-RADIOv2 does not reach a breaking point, it comes very close to one, which we could attribute to its mixed supervision weakening the reliability of part-level representations.
In contrast, TIPS and VGGT exhibit clear breakdowns, indicating that their learned features are less robust when faced with unseen viewpoints. For TIPS, this instability likely stems from its focus on semantic alignment rather than spatial consistency. VGGT breaks for a different reason: it is designed for multi-view fusion, and receiving only a single reference view places it outside its intended use. Relative to the baseline, where validation bins are also included in memory, these observations confirm that the sudden declines arise from missing reference views rather than instability in the evaluation pipeline. Overall, the results show that self-supervised encoders such as the DINO-based models produce more geometry-aware features under strong viewpoint shifts, while multimodal or distilled approaches are more brittle. DINOv2’s stability makes it the most reliable choice for applications that must handle abrupt perspective changes.
@@ 4.3 Experiment C: memory size robustness
In this final experiment, we analyze how memory bank size affects viewpoint generalization, comparing the performance gains across difficulty levels against the added computational cost. We follow the Hummingbird paper [hummingbird] and vary the number of entries in the memory bank. Three memory sizes are evaluated: 320k, 640k, and 1,024k. For each size, we report segmentation accuracy and measure computational cost, including runtime and resource usage.
@@ 4.3.1 Results
The results in Table 4 show that increasing memory generally improves mIoU, with the largest gains observed under the Easy difficulty. Doubling memory from 320k to 640k increases performance by about 0.030 mIoU on Easy, while a further increase to 1,024k yields a smaller improvement of 0.017 (see Table 10 in Appendix E.4). DINOv3 achieves the highest absolute performance across all memory sizes, followed by DINO and DINOv2. Weaker models such as C-RADIOv2, SigLIP2, TIPS and VGGT show the greatest relative gains from increasing the memory. All but the former improve by more than 0.048 when moving from 320k to 1,024k, with most of that improvement coming from the first memory doubling.
The runtime analysis in Table 11 in Appendix E.4 shows that larger memory banks increase computational cost. The highest configuration adds about two hours per job. Total memory usage remains similar (189–214 GB), but CPU efficiency declines as memory grows: for example, C-RADIOv2 drops from 16.35% at 320k to 12.52% at 1,024k.
Table 4: Comparison of memory-bank size on segmentation performance. Mean mIoU scores and standard deviations are shown for three memory sizes ((a) 320k, (b) 640k, (c) 1,024k) and four difficulty levels. Bold values mark the best score per difficulty-memory configuration. Increasing memory generally improves performance, especially for weaker encoders such as SigLIP2. The DINO models remain the top performers across all settings, but have minimal gains past 640k.
@@ 5 Qualitative analysis of segmentations
In addition to quantitative evaluations, we conducted a qualitative inspection of predicted segmentation masks to better understand model behaviour under the large-angle changes of the Extreme difficulty. The illustrations can be found in Appendix E.5.
@@ 5.4 Failures
Failure cases, such as the toy dragon shown in Figure 30, typically arise when the model retrieves visually similar but geometrically inconsistent patches. This may lead to boundary inflation. Such failures explain the sharp performance drops observed in certain models (e.g., TIPS) at larger angular deviations, where cross-view consistency becomes more challenging.
@@ 6 Conclusion
We studied the ability of frozen ViT encoders to generalize across unseen camera viewpoints in an in-context segmentation setting, using the Hummingbird architecture and MVImgNet dataset. Our benchmark covered three aspects: controlled evaluation under view shifts, robustness through breaking point analysis, and the effect of memory bank size. Our findings show that self-supervised encoders, particularly DINO, DINOv2, and DINOv3, provide more stable geometry-aware features under viewpoint changes, while multimodal and distilled approaches are more brittle. VGGT underperforms because its multi-view architecture and training objective are likely misaligned with the single-view, patch-based retrieval setup used in our experiments. Larger memory banks help weaker models but offer diminishing returns for stronger encoders, raising questions about the cost-benefit trade-off.
Taken together, these results highlight the strengths and limitations of current ViTs for multi-view perception. While modern self-supervised training yields robust features, none of the models are immune to viewpoint-induced degradation. Addressing this gap will likely require objectives or architectures that enforce 3D consistency. Future work could include extending this evaluation to multi-object or multi-class scenes, wider angular ranges, and compound rotations, providing a broader testbed for building truly viewpoint-robust representations.
@@ Appendix A Licenses for assets
VGGT is released under it’s own custom VGGT License (v1 Last Updated: July 29, 2025). Non-commercial use is permitted, allowing academic and research use.
@@ Appendix B The MVImgNet dataset
The statistics for the reorganized subset of MVImgNet [yu2023mvimgnetlargescaledatasetmultiview] we used are shown in Table 6. The selected subset includes 15 object classes, each represented across 7 standardized angle bins: [0, 15, 30, 45, 60, 75, 90]. Figure 4 shows an instance of each angle for each object class.
Table 6: Angle selection accuracy per class. Mean and standard deviation of angular error (in degrees) for each object class after selecting views closest to the predefined angle bins. Bin image counts marked with $\dagger$ indicate the bin sizes used for the models DINOv3 and VGGT. Due to a misconfiguration, the dataset sizes were reduced for eight classes for these two models.
@@
99 Coat rack 1.41 -0.02 97 (9)
100 Guitar stand 1.53 0.04 218 (3)
113 Ceiling lamp 1.64 -0.03 154 (75)
125 Toilet 1.31 0.02 58 (5)
126 Sink 1.20 -0.12 30 (5)
152 Strings 1.25 0.03 192 (121)
@@
Figure 4: Multi-view categories. One representative instance is shown for each of the 15 selected MVImgNet classes across all viewpoint bins (0–90). Based on the category taxonomy of MVImgNet [yu2023mvimgnetlargescaledatasetmultiview], we have examples from both the food and artifacts parents class, which encompass the following children classes: fruits and vegetables, toy dolls and a series of instruments: furniture, kitchen ware, home appliances, and entertainment.
@@ Appendix C Model specifics
VGGT Geometry-Grounded 504 14 1296 1024 4
The models differ in their native input resolutions: $224\times 224$ for CLIP ViT-B/16 [CLIP], DINO ViT-B/16 [DINO], DINOv3 ViT-B/16 [DINOv3], VGGT [VGGT] and C-RADIOv2 ViT-B/16-CPE [radiov2.5, am-radio]; $512\times 512$ for DINOv2 ViT-B/14 [DINOv2]; $504\times 504$ for SigLIP2 B/16-512 [SIGLIPv2]; and $448\times 448$ for TIPS ViT-B/14-HR [tips_paper].
VGGT employs the ViT-L backbone used in DINOv2, consisting of 24 transformer blocks that alternate between frame-wise and global self-attention. Each input image is patchified into tokens using a pretrained DINOv2 [DINOv2] tokenizer, maintaining compatibility with other ViT-based encoders in our study. The alternating-attention design allows the encoder to integrate information across frames while refining features within each image. Tokens from selected intermediate blocks are later used by a DPT [DPT] head for upsampling. Compared to the standard ViT-B encoders (e.g., DINO, CLIP, DINOv2), VGGT’s larger capacity and dual-attention structure make it architecturally distinct.
@@ Appendix D Computational resources
Table 8 outlines the specific hardware and software configurations used to execute all experiments, ensuring reproducibility and accounting for memory constraints. All experiments were executed on the Snellius HPC cluster.
Table 8: Experimental hardware and configuration.
Category Component/value Notes
Hardware 4 $\times$ A100 GPUs High-performance cluster
72 CPU cores
Max wall time 30 hours per job
Software/settings Search index distribution FAISS sharding across GPUs
Batch size 4
Random seed Python, NumPy, PyTorch, CUDA (set to 42)
Appendix E Additional results
E.1 Reproduction
Table 9: Reproduction results. We report accuracy (%) at different evaluation scales. “Reported” indicates values taken from prior published results, while “Reproduced” refers to our reproduction with a batch size (BS). Bold values highlight the best result for each model-scale configuration. Overall, our reproduced scores are close to or exceed the reported results in the original paper, with reproduced results differing by less than 3%.

### v2

@@ 4.1.2 Discussion
Although none of the models were explicitly trained for viewpoint understanding, the results show clear differences in their ability to generalize across views. As seen in the previous section, DINOv3 performs best, followed by DINOv2 and DINO. This suggests that the features of the DINO models, compared to other models, are more stable under viewpoint changes and better support shape-related cues. This consistency may arise from their self-supervised training objective, which explicitly encourages feature alignment across multiple augmentations of the same image and more stable relationships between local and global representations. DINOv3 further benefits from large-scale pre-training and the Gram Anchoring mechanism [27], which might contribute to more stable local features and stronger geometric consistency under viewpoint changes.
While DINO performed better under the easier difficulties, where multiple views of the object were available, DINOv2 obtained a higher mIoU under the Extreme difficulty with only 1 reference bin. One possible interpretation is that DINO’s simpler self-distillation loss favors interpolation across views, whereas DINOv2’s larger-scale self-supervised pre-training better supports generalization to more extreme viewpoints. CLIP generally ranked third, reflecting some ability to encode visual-semantic regularities despite being trained for image-text alignment rather than geometric consistency. By comparison, SigLIP2, TIPS, and C-RADIOv2 performed substantially worse, suggesting that objectives centered on semantic matching or mixed supervision may weaken part-level consistency needed for viewpoint-robust segmentation.
Figure 2: Segmentation performance across viewpoint bins. Each subplot (a)–(d) corresponds to a different difficulty level. Bars show mIoU scores on unseen validation bins, while dots show performance on the reference bins stored in the memory bank. The figure illustrates how segmentation accuracy declines as the validation angles move farther from the reference viewpoints. DINOv3, DINOv2, DINO, and CLIP maintain higher, more stable performance as angular distance increases, whereas other models degrade more quickly. Overall, the trend highlights how self-supervised pre-training leads to smoother cross-view generalization.
@@ 4.2 Experiment B: breaking point analysis
In this experiment, we analyze whether the models experience a sudden failure in viewpoint generalization, and if so, at what angle. The goal is to identify the angular range within which performance remains stable under the Extreme difficulty. In particular, we perform a breaking point analysis under the Extreme setting from experiment A, where only a single reference bin (0) is available in the memory and models are evaluated across the remaining bins. We define the breaking point as the earliest validation bin where a model’s normalized mIoU score $\tilde{m}_{i}$ drops significantly compared to the previous bin ($\tilde{m}_{i-1}$). Specifically, we compute the performance drop $\Delta_{i}$ per bin $i$ as
$\Delta_{i}$ $=\tilde{m}_{i}-\tilde{m}_{i-1},$ (2)
$\tilde{m}_{i}$ $=\frac{m_{i}}{m_{0^{\circ}}},$ (3)
where the normalization by the 0 bin (Equation 3) ensures that performance drops reflect viewpoint sensitivity rather than overall model scale or performance.
@@ 4.2.1 Results
As shown in Figure 3, all models exhibit a monotonic decrease in normalized mIoU as the validation viewpoint angle increases. The drop is steepest at small viewpoint changes and gradually flattens at larger angles, indicating that most performance degradation occurs at the initial viewpoint shifts.
TIPS is the only model that reaches a breaking point, occurring at the $30^{\circ}$ bin (see Table 3), with a normalized mIoU drop of 0.1148-0.1148. C-RADIOv2 shows a relatively sharp early decline but remains below the breaking-point threshold. DINOv3, DINOv2, and CLIP exhibit the smoothest degradation, followed by DINO and SigLIP2.
Figure 3: Normalized mIoU under viewpoint shifts. Each curve shows the normalized mIoU of a model relative to its performance in the $0^{\circ}$ reference bin. The horizontal axis corresponds to increasing viewpoint angles (validation bins). All models exhibit a monotonic decrease with diminishing slope, indicating that the steepest performance degradation occurs at the initial viewpoint shifts, followed by progressively flatter declines at larger angles. DINOv3, DINOv2, and CLIP degrade smoothly, followed by DINO and SigLIP2, whereas TIPS and C-RADIOv2 display sharper early drops, suggesting weaker robustness to small viewpoint changes.
Table 3: Breaking points. For each model, we report the breaking point validation bin where the normalized mIoU drop exceeds 10% ($\Delta_{i}\leq-0.1$) relative to the previous bin. “None” indicates gradual degradation. Only TIPS reaches a breaking point. C-RADIOv2 is close to having a breaking point, but it doesn’t pass the threshold as per our definition. DINO, DINOv2, CLIP, and SigLIP2 maintain smoother decline curves. The biggest drops for all models occur between 15 and 30 as seen in Figure 3.
Model Breaking point bin Biggest drop
CLIP ViT-B/16 None 0.0267-0.0267
DINO ViT-B/16 None 0.0559-0.0559
DINOv2 ViT-B/14 None 0.0273-0.0273
DINOv3 ViT-B/16 None 0.0214-0.0214
C-RADIOv2 ViT-B/16-CPE None 0.0969-0.0969
SigLIP2 B/16-512 None 0.0550-0.0550
@@ 4.2.2 Discussion
The breaking point analysis highlights differences in robustness across models and shows how their pre-training objectives affect stability under viewpoint shifts. DINOv2 and DINOv3 are the most stable, maintaining gradual degradation without sudden drops. CLIP shows a similar trend, with slightly weaker overall performance. The normalized performances of DINO and SigLIP2 also decrease steadily, with a slight increase at the 90 bin. Interestingly, CLIP and SigLIP2 preserve some resilience despite being trained for image-text alignment rather than spatial structure. Although C-RADIOv2 does not reach a breaking point, it comes very close to one, which we could attribute to its mixed supervision weakening the reliability of part-level representations. In contrast, TIPS exhibits a clear breakdown, suggesting that its learned features are less robust when faced with unseen viewpoints. This instability may stem from its focus on semantic alignment rather than spatial consistency. Relative to the baseline that includes validation bins in the memory, the observed drops can be attributed to missing reference views rather than to instability in the evaluation pipeline. Overall, the results suggest that self-supervised encoders such as the DINO-based models produce more geometry-aware features under strong viewpoint shifts, while multimodal or distilled approaches may be more brittle. DINOv2 and DINOv3’s stability make them the most reliable choice for applications that must handle abrupt perspective changes.
@@ 4.3 Experiment C: memory size robustness
In this final experiment, we analyze how memory bank size affects viewpoint generalization, comparing the performance gains across difficulty levels against the added computational cost. We follow the Hummingbird paper [1] and vary the number of entries in the memory bank. Three memory sizes are evaluated: 320k, 640k, and 1,024k. For each size, we report the segmentation accuracy.
@@ 4.3.1 Results
The results in Table 4 show that increasing the memory bank generally improves mIoU, with the largest gains observed under the Easy difficulty. Doubling the memory from 320k to 640k increases performance by approximately 0.030 mIoU on Easy, while a further increase to 1,024k yields a smaller improvement of 0.017 (see Table 9 in Appendix F.3).
DINOv3 achieves the highest absolute performance across all memory sizes, followed by DINO and DINOv2. Weaker models such as C-RADIOv2, SigLIP2, and TIPS benefit more from additional memory, improving by more than 0.048 mIoU when increasing the memory from 320k to 1,024k, with most of the gain coming from the first doubling of memory.
Increasing the memory bank introduces additional practical considerations, as documented in Appendix F.3.
Table 4: Comparison of memory-bank size on segmentation performance. Mean mIoU scores and standard deviations are shown for 3 memory sizes ((a) 320k, (b) 640k, (c) 1,024k) and 4 difficulty levels. Bold values mark the best score per difficulty-memory configuration. Increasing memory generally improves performance, especially for weaker encoders such as SigLIP2. The DINO models remain the top performers across all settings, but have minimal gains past 640k.
@@ 5 Qualitative analysis of segmentations
In addition to quantitative evaluations, we conducted a qualitative inspection of predicted segmentation masks to better understand model behavior under the large-angle changes of the Extreme difficulty. The illustrations mentioned in this section can be found in Appendix G.
@@ 5.4 Failures
Failure cases, such as the toy dragon shown in Figure 30, typically arise when the model retrieves visually similar but geometrically inconsistent patches. This may lead to boundary inflation. Such failures may explain the sharp performance drops observed in certain models (e.g., TIPS) at larger angular deviations, where cross-view consistency becomes more challenging.
@@ 6 Conclusion
We studied the ability of frozen ViT encoders to generalize across unseen camera viewpoints in an in-context segmentation setting, using the Hummingbird architecture and MVImgNet dataset. Our benchmark covered three aspects: controlled evaluation under view shifts, robustness through breaking point analysis, and the effect of memory bank size. Our findings show that self-supervised encoders, particularly DINO, DINOv2, and DINOv3, achieve higher performance under viewpoint changes than multimodal and mixed-supervision approaches. Increasing the memory bank size benefits weaker models and offers diminishing returns for stronger encoders.
Taken together, these results highlight the strengths and limitations of current ViTs for multi-view perception. While self-supervised training objectives are associated with improved robustness, none of the models are immune to viewpoint-induced degradation. Addressing this gap will likely require objectives or architectures that more explicitly account for 3D or viewpoint consistency. Future work could include extending this evaluation to multi-object or multi-class scenes, wider angular ranges, and compound rotations, providing a broader testbed for building truly viewpoint-robust representations.
@@ Appendix B The MVImgNet dataset
The statistics for the reorganized subset of MVImgNet [34] we used are shown in Table 6. The selected subset includes 15 object classes, each represented across 7 standardized angle bins: 0, 15, 30, 45, 60, 75, and 90. Figure 4 shows an instance of each angle for each object class.
Table 6: Angle selection accuracy per class. Mean and standard deviation of angular error (in degrees) for each object class after selecting views closest to the predefined angle bins.
@@
99 Coat rack 1.41 -0.02 97
100 Guitar stand 1.53 0.04 218
113 Ceiling lamp 1.64 -0.03 154
125 Toilet 1.31 0.02 58
126 Sink 1.20 -0.12 30
152 Strings 1.25 0.03 192
@@
Figure 4: Multi-view categories. A representative instance is shown for each of the 15 selected MVImgNet classes across all viewpoint bins (0–90). Based on the category taxonomy of MVImgNet [34], we have examples from both the food and artifacts parent classes, which encompass the following child classes: fruits and vegetables, toy dolls, and a series of instruments: furniture, kitchen ware, home appliances, and entertainment.
@@ Appendix C Model specifics
The models differ in their native input resolutions: $224\times 224$ for CLIP ViT-B/16 [18], DINO ViT-B/16 [3], DINOv3 ViT-B/16 [27] and C-RADIOv2 ViT-B/16-CPE [8, 19]; $512\times 512$ for DINOv2 ViT-B/14 [16]; $504\times 504$ for SigLIP2 B/16-512 [29]; and $448\times 448$ for TIPS ViT-B/14-HR [13].
@@ Appendix D Computational resources
All experiments were executed on the Snellius HPC cluster to ensure high-performance processing and reproducibility. We ran our experiments on a single A100 node consisting of 4 $\times$ NVIDIA A100 GPUs and 72 CPU cores, with a maximum wall time of 30 hours allocated per job to account for memory and processing constraints.
Regarding software and environment settings, the search index was distributed using FAISS sharding across all available GPUs. We utilized a batch size of 4 for all runs. To ensure experimental consistency, a global random seed of 42 was set across Python, NumPy, PyTorch, and CUDA.
Appendix E Reproduction results
Table 8: Reproduction results. We report accuracy (%) at different evaluation scales. “Reported” indicates values taken from prior published results, while “Reproduced” refers to our reproduction with a batch size (BS). Bold values highlight the best result for each model-scale configuration. Overall, our reproduced scores are close to or exceed the reported results in the original paper.